OpenStudy (anonymous):

Solve the equation: |x-3|=8

jimthompson5910 (jim_thompson5910):

Hint: if |x| = k, then x = k or x = -k for some positive number k

jimthompson5910 (jim_thompson5910):

So if |x-3|=8 then x-3 = 8 or x-3 = -8

OpenStudy (anonymous):

Your seriously a genius! Thanks :)

jimthompson5910 (jim_thompson5910):

lol sure thing, glad to be of help